OpenStudy (anonymous):

Solve 15x = 3y - 27 for y. Put your answer in the form y=mx+b with no spaces

OpenStudy (anonymous):

put 3y alone on one side: 3y - 27 = 15x +27 3y = 15x + 27 isolate the y: /3 y=5x+9 Please pick me as best answer:)